OpenStudy (anonymous):

Ok one more 16s^6-2g^6 so 2s(2s)^2-2g(g)^2

OpenStudy (anonymous):

hmmm....is that your simplification?

OpenStudy (anonymous):

Yes am I wrong is this not the answer?

OpenStudy (anonymous):

yeah you have to notice that (2s)^2 = 4s^2 and so multiplying that by 2s would be 8s^3 and 2g*g^2 = 2g^3

OpenStudy (anonymous):

Then am I to assume that (2s)^4 (2g)3

OpenStudy (anonymous):

no because then you have 16s^4

OpenStudy (anonymous):

one way to look at it is seperating the equation into a product of binomials

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on let me write something up first on paper to see if its correct

OpenStudy (anonymous):

ok i got something

OpenStudy (anonymous):

2s(2s)^5-2g(g)^5

OpenStudy (anonymous):

so notice factoring out a 2 you get \[2( 8s ^{6} - g^{6})\] and doing a little algebra we get \[2( (2s^{2})^{3} - (g^{2})^{3})\] and so using the difference of cubes we get \[2[ (2s^{2} - g^{2})* ( (2s^{2})^{2} +2s^{2}g^{2} + (g^{2})^{2})\]

OpenStudy (anonymous):

what you did is incorrect since (2s)^5 = 32s^5 and multiplying by 2s would be 64s^6

OpenStudy (anonymous):

How about: \[-2(g^2-2s^2)(g^4+2g^2 s^2+4s^4)\]

OpenStudy (anonymous):

dude if you look at what i wrote what you wrote is the exact same

OpenStudy (anonymous):

Yes, after removing the open bracket symbol they are equivalent. Good work robotsandtoast.